# PASCAL: Precise and Efficient ANN- SNN Conversion using Spike Accumulation and Adaptive Layerwise Activation

### v1

@@ A.2 Effect of threshold of Van Der Eijk’s A on the Final Metric
Figure 3: Layerwise AL-metric for different models and datasets, with the threshold $\alpha=\frac{k}{|\mathbf{L_{tot}}|}$ for Agreement, where $\mathbf{L_{tot}}$ is the number of layers, computed using VGG-16 for CIFAR-100. (a) $k=1$ (b) $k=0.75$ (c) $k=0.5$
@@ A.3 Introduction of Datasets
CIFAR-10. The CIFAR-10 dataset (Krizhevsky et al., 2009) consists of 60000 $32\times 32$ images in 10 classes. There are 50000 training images and 10000 test images.
CIFAR-100. The CIFAR-100 dataset (Krizhevsky et al., 2009) consists of 60000 $32\times 32$ images in 100 classes. There are 50000 training images and 10000 test images.
ImageNet. We use the ILSVRC 2012 dataset (Russakovsky et al., 2015), which consists 1,281,167 training images and 50000 testing images.
@@ A.4 Comparison of the number of MACs and ACs between ANN and SNN
The table below shows the number of operations for both convolution and fully connected layers. Table 8 shows the notation used for this analysis. The energy costs of addition and multiplication in 45nm CMOS is shown in Table 10.
Table 8: Notation used in computing MACs and ACs
@@
Table 10: Energy costs of addition and multiplication in 45nm CMOS (Horowitz, 2014)
@@ A.5 Layerwise count of operations in the VGG-16 architecture for CIFAR-10 and CIFAR-100
Conv1_1 (Convolution) $32\times 32\times 3$ $32\times 32\times 64$ $32\times 32\times 3\times 3\times 3\times 64=1,769,472$
Conv1_2 (Convolution) $32\times 32\times 64$ $32\times 32\times 64$ $32\times 32\times 64\times 3\times 3\times 64=37,748,736$
@@
Conv2_1 (Convolution) $16\times 16\times 64$ $16\times 16\times 128$ $16\times 16\times 64\times 3\times 3\times 128=18,874,368$
Conv2_2 (Convolution) $16\times 16\times 128$ $16\times 16\times 128$ $16\times 16\times 128\times 3\times 3\times 128=37,748,736$
@@
Conv3_1 (Convolution) $8\times 8\times 128$ $8\times 8\times 256$ $8\times 8\times 128\times 3\times 3\times 256=18,874,368$
Conv3_2 (Convolution) $8\times 8\times 256$ $8\times 8\times 256$ $8\times 8\times 256\times 3\times 3\times 256=37,748,736$
Conv3_3 (Convolution) $8\times 8\times 256$ $8\times 8\times 256$ $8\times 8\times 256\times 3\times 3\times 256=37,748,736$
@@
Conv4_1 (Convolution) $4\times 4\times 256$ $4\times 4\times 512$ $4\times 4\times 256\times 3\times 3\times 512=18,874,368$
Conv4_2 (Convolution) $4\times 4\times 512$ $4\times 4\times 512$ $4\times 4\times 512\times 3\times 3\times 512=37,748,736$
Conv4_3 (Convolution) $4\times 4\times 512$ $4\times 4\times 512$ $4\times 4\times 512\times 3\times 3\times 512=37,748,736$
@@
Conv5_1 (Convolution) $2\times 2\times 512$ $2\times 2\times 512$ $2\times 2\times 512\times 3\times 3\times 512=9,437,184$
Conv5_2 (Convolution) $2\times 2\times 512$ $2\times 2\times 512$ $2\times 2\times 512\times 3\times 3\times 512=9,437,184$
Conv5_3 (Convolution) $2\times 2\times 512$ $2\times 2\times 512$ $2\times 2\times 512\times 3\times 3\times 512=9,437,184$
@@ A.6 Layerwise count of operations in the VGG-16 architecture for ImageNet
Conv1_1 (Convolution) $224\times 224\times 3$ $224\times 224\times 64$ $224\times 224\times 3\times 3\times 3\times 64=86,704,128$
Conv1_2 (Convolution) $224\times 224\times 64$ $224\times 224\times 64$ $224\times 224\times 64\times 3\times 3\times 64=1,849,688,064$
@@
Conv2_1 (Convolution) $112\times 112\times 64$ $112\times 112\times 128$ $112\times 112\times 64\times 3\times 3\times 128=924,844,032$
Conv2_2 (Convolution) $112\times 112\times 128$ $112\times 112\times 128$ $112\times 112\times 128\times 3\times 3\times 128=1,849,688,064$
@@
Conv3_1 (Convolution) $56\times 56\times 128$ $56\times 56\times 256$ $56\times 56\times 128\times 3\times 3\times 256=924,844,032$
Conv3_2 (Convolution) $56\times 56\times 256$ $56\times 56\times 256$ $56\times 56\times 256\times 3\times 3\times 256=1,849,688,064$
Conv3_3 (Convolution) $56\times 56\times 256$ $56\times 56\times 256$ $56\times 56\times 256\times 3\times 3\times 256=1,849,688,064$
@@
Conv4_1 (Convolution) $28\times 28\times 256$ $28\times 28\times 512$ $28\times 28\times 256\times 3\times 3\times 512=924,844,032$
Conv4_2 (Convolution) $28\times 28\times 512$ $28\times 28\times 512$ $28\times 28\times 512\times 3\times 3\times 512=1,849,688,064$
Conv4_3 (Convolution) $28\times 28\times 512$ $28\times 28\times 512$ $28\times 28\times 512\times 3\times 3\times 512=1,849,688,064$
@@
Conv5_1 (Convolution) $14\times 14\times 512$ $14\times 14\times 512$ $14\times 14\times 512\times 3\times 3\times 512=462,422,016$
Conv5_2 (Convolution) $14\times 14\times 512$ $14\times 14\times 512$ $14\times 14\times 512\times 3\times 3\times 512=462,422,016$
Conv5_3 (Convolution) $14\times 14\times 512$ $14\times 14\times 512$ $14\times 14\times 512\times 3\times 3\times 512=462,422,016$
@@ A.8 Training Configurations
In order to implement PASCAL, we adopt the code framework of QCFS  (Bu et al., 2023). We use the Stochastic Gradient Descent optimizer (Bottou, 2012) with a momentum parameter of 0.9. The initial learning rate is set to 0.1 for CIFAR-10 and ImageNet, and 0.02 for CIFAR-100. A cosine decay scheduler (Loshchilov & Hutter, 2016) is used to adjust the learning rate. We apply a $5\times 10^{-4}$ weight decay for CIFAR datasets while applying a $1\times 10^{-4}$ weight decay for ImageNet.
@@ A.9 Layerwise spike rate over the test set
We plot the spike rate over the entire test set for various model architectures and datasets in Figure 4
@@
Figure 4: Layerwise Spike Rate over the test-set, (a) for $L=4$ using VGG-16 for both CIFAR-10 and CIFAR-100, (b) for $L=8$ using ResNet-18 for both CIFAR-10 and CIFAR-100, and (c) for $L=16$ using VGG-16 for ImageNet
A.10 Quantifying the overhead due to the spike accumulator
We first note that the IF layer acts only on the output dimensions. Therefore, the number of operations is considerably smaller than MatMul layers, which involve operations proportional to the product of the input channels and output channels. We perform the analysis for convolutions first.
A.10.1 Convolution
We assume constant quantization step $L_{n}$ for further calculations in this section. As per Algorithm 1, we know that number of soft reset operations is $13L_{n}-$ 1. Also, the number of spike accumulation operations, i.e. addition and subtraction of the spike counter, is $12L_{n}-$ 1. In total, each IF layer performs $(5L_{n}-2)\times C_{0}\times H_{0}\times W_{0}$ accumulation operations. Therefore, we compute the ratio between the number of operations in the IF layer to the number of operations in the previous convolution layer is
#$\frac{\#OP_{IF}}{\#OP_{Conv}}\approx\frac{5L_{n}-1}{C_{i}\cdot K_{h}\cdot K_{w}\cdot SpikeRate_{l}}$ (35)
We refer to an extension of (Horowitz, 2014)’s energy calculations in the table below.
Table 14: Energy costs of Integer addition and multiplication in 45nm CMOS (Horowitz, 2014)
INT ADD (8 bit) $0.03pJ$
INT ADD (32 bit) $0.1pJ$
We note that the spike accumulation operations can be reduced to 8-bit addition and subtraction, since we use relatively lower values of $L_{n}$ in this work. Therefore, the cost of $12L_{n}-$ 1 operations would reduce even further, becoming $\frac{0.03}{0.9}=\frac{1}{30}^{th}$ of a floating point addition.
After these considerations, the final ratio is
#$\frac{\#OP_{IF}}{\#OP_{Conv}}=\frac{(2L_{n}-1)\frac{1}{30}+(3L_{n}-1)}{C_{i}\cdot K_{h}\cdot K_{w}\cdot SpikeRate_{l}}$ (36)
As an illustration, we choose $K_{h}=K_{w}=3$ and $C_{i}=256$. This is true for a typical layer in the middle of a network. For $L_{n}=8$ and $SpikeRate_{l}=0.75$, a value commonly used, we get the ratio to be 0.016. In other words, the operations in an IF layer amount to $\approx 1.5\%$ of the computation in the previous convolution layer. Hence, we have neglected this calculation from the main paper.
A.10.2 Fully Connected (FC) Layer
Following a similar reasoning to the previous section, we can deduce that
#$\frac{\#OP_{IF}}{\#OP_{FC}}=\frac{(2L_{n}-1)\frac{1}{30}+(3L_{n}-1)}{C_{i}\cdot SpikeRate_{l}}$ (37)

### v2

@@ A.2 Effect of threshold of Van Der Eijk’s A on the Final Metric
Figure 3: Layerwise AL-metric with the threshold $\alpha=\frac{k}{|\mathbf{L_{tot}}|}$ for Agreement, where $\mathbf{L_{tot}}$ is the number of layers, computed using VGG-16 for CIFAR-100. (a) $k=1$ (b) $k=0.75$ (c) $k=0.5$.
@@ A.3 Introduction of Datasets
CIFAR-10. The CIFAR-10 dataset (krizhevsky2009learning) consists of 60000 $32\times 32$ images in 10 classes. There are 50000 training images and 10000 test images.
CIFAR-100. The CIFAR-100 dataset (krizhevsky2009learning) consists of 60000 $32\times 32$ images in 100 classes. There are 50000 training images and 10000 test images.
ImageNet. We use the ILSVRC 2012 dataset (russakovsky2015imagenet), which consists of 1,281,167 training images and 50000 testing images.
@@ A.4 Comparison of the number of MACs and ACs between ANN and SNN
The table below shows the number of operations for both convolution and fully connected layers. Table 8 shows the notation used for this analysis. The energy cost of addition and multiplication in 45nm CMOS is shown in Table 10.
Table 8: Notation used in computing MACs and ACs.
@@
Table 10: Energy cost of addition and multiplication in 45nm CMOS (horowitz20141).
@@ A.5 Layerwise count of operations in the VGG-16 architecture for CIFAR-10 and CIFAR-100
Conv1_1 (Convolution) $\times 32\times 3 32×32$$\times 32\times 64$ $32\times 32\times 3\times 3\times 3\times 64=1,769,472$
Conv1_2 (Convolution) $\times 32\times 64 32×32$$\times 32\times 64$ $32\times 32\times 64\times 3\times 3\times 64=37,748,736$
@@
Conv2_1 (Convolution) $\times 16\times 64 16×16$$\times 16\times 128$ $16\times 16\times 64\times 3\times 3\times 128=18,874,368$
Conv2_2 (Convolution) $\times 16\times 128 16×16$$\times 16\times 128$ $16\times 16\times 128\times 3\times 3\times 128=37,748,736$
@@
Conv3_1 (Convolution) $\times 8\times 128 8×8$$\times 8\times 256$ $8\times 8\times 128\times 3\times 3\times 256=18,874,368$
Conv3_2 (Convolution) $\times 8\times 256 8×8$$\times 8\times 256$ $8\times 8\times 256\times 3\times 3\times 256=37,748,736$
Conv3_3 (Convolution) $\times 8\times 256 8×8$$\times 8\times 256$ $8\times 8\times 256\times 3\times 3\times 256=37,748,736$
@@
Conv4_1 (Convolution) $\times 4\times 256 4×4$$\times 4\times 512$ $4\times 4\times 256\times 3\times 3\times 512=18,874,368$
Conv4_2 (Convolution) $\times 4\times 512 4×4$$\times 4\times 512$ $4\times 4\times 512\times 3\times 3\times 512=37,748,736$
Conv4_3 (Convolution) $\times 4\times 512 4×4$$\times 4\times 512$ $4\times 4\times 512\times 3\times 3\times 512=37,748,736$
@@
Conv5_1 (Convolution) $\times 2\times 512 2×2$$\times 2\times 512$ $2\times 2\times 512\times 3\times 3\times 512=9,437,184$
Conv5_2 (Convolution) $\times 2\times 512 2×2$$\times 2\times 512$ $2\times 2\times 512\times 3\times 3\times 512=9,437,184$
Conv5_3 (Convolution) $\times 2\times 512 2×2$$\times 2\times 512$ $2\times 2\times 512\times 3\times 3\times 512=9,437,184$
@@ A.6 Layerwise count of operations in the VGG-16 architecture for ImageNet
Conv1_1 (Convolution) $\times 224\times 3 224×224$$\times 224\times 64$ $224\times 224\times 3\times 3\times 3\times 64=86,704,128$
Conv1_2 (Convolution) $\times 224\times 64 224×224$$\times 224\times 64$ $224\times 224\times 64\times 3\times 3\times 64=1,849,688,064$
@@
Conv2_1 (Convolution) $\times 112\times 64 112×112$$\times 112\times 128$ $112\times 112\times 64\times 3\times 3\times 128=924,844,032$
Conv2_2 (Convolution) $\times 112\times 128 112×112$$\times 112\times 128$ $112\times 112\times 128\times 3\times 3\times 128=1,849,688,064$
@@
Conv3_1 (Convolution) $\times 56\times 128 56×56$$\times 56\times 256$ $56\times 56\times 128\times 3\times 3\times 256=924,844,032$
Conv3_2 (Convolution) $\times 56\times 256 56×56$$\times 56\times 256$ $56\times 56\times 256\times 3\times 3\times 256=1,849,688,064$
Conv3_3 (Convolution) $\times 56\times 256 56×56$$\times 56\times 256$ $56\times 56\times 256\times 3\times 3\times 256=1,849,688,064$
@@
Conv4_1 (Convolution) $\times 28\times 256 28×28$$\times 28\times 512$ $28\times 28\times 256\times 3\times 3\times 512=924,844,032$
Conv4_2 (Convolution) $\times 28\times 512 28×28$$\times 28\times 512$ $28\times 28\times 512\times 3\times 3\times 512=1,849,688,064$
Conv4_3 (Convolution) $\times 28\times 512 28×28$$\times 28\times 512$ $28\times 28\times 512\times 3\times 3\times 512=1,849,688,064$
@@
Conv5_1 (Convolution) $\times 14\times 512 14×14$$\times 14\times 512$ $14\times 14\times 512\times 3\times 3\times 512=462,422,016$
Conv5_2 (Convolution) $\times 14\times 512 14×14$$\times 14\times 512$ $14\times 14\times 512\times 3\times 3\times 512=462,422,016$
Conv5_3 (Convolution) $\times 14\times 512 14×14$$\times 14\times 512$ $14\times 14\times 512\times 3\times 3\times 512=462,422,016$
@@ A.8 Training Configurations
In order to implement PASCAL, we adopt the code framework of QCFS  bu2023optimal. We use the Stochastic Gradient Descent optimizer (bottou2012stochastic) with a momentum parameter of 0.9. The initial learning rate is set to 0.1 for CIFAR-10 and ImageNet, and 0.02 for CIFAR-100. A cosine decay scheduler (loshchilov2016sgdr) is used to adjust the learning rate. We apply a $5\times 10^{-4}$ weight decay for CIFAR datasets while applying a $1\times 10^{-4}$ weight decay for ImageNet.
@@ A.9 Layerwise spike rate over the test set
We plot the spike rate over the entire test set for various model architectures and datasets in Figure 4.
@@
Figure 4: Layerwise Spike Rate over the test-set, (a) for $L=4$ using VGG-16 for both CIFAR-10 and CIFAR-100, (b) for $L=8$ using ResNet-18 for both CIFAR-10 and CIFAR-100, and (c) for $L=16$ using VGG-16 for ImageNet.
A.10 Hardware Quantification
A.10.1 Energy
We implemented a custom SNN accelerator in-house to quantify the energy consumed per inference with and without the PASC IF layer. The accelerator is a modified version of the LoAS accelerator (yin2024loas), enhanced to support the PASC IF layer. We report the energy metrics for the VGG-16 SNN, over the CIFAR-10 test set.
Table 14: Energy metrics of VGG-16 SNN with and without PASC IF.
Type of IF Total Inference Energy Energy Contributed by the IF Layer
Default IF (T = 4) 445.47 mJ 0.08 mJ
PASC IF (T = 4) 446.02 mJ 0.63 mJ
Our results show that the VGG-16 SNN with PASC IF neurons consumes only 0.1% more energy compared to the default IF implementation.
In an effort to further validate the energy efficiency results, we re-implemented an ANN accelerator SparTen in addition to the aforementioned SNN accelerator LoAS  yin2024loas. Both the accelerators were designed to support INT8 quantization for the weights. In addition, SparTen was enhanced to use the QCFS function to implement quantized activation. We synthesized the Register Transfer Level (Verilog RTL) design of the accelerators using the Synopsys Design Compiler on 40 nm CMOS technology to determine the operating frequency (560 MHz) and power consumption. We then developed a cycle-accurate simulator of the respective accelerators to estimate inference latency and energy. Table  15 summarizes the inference energy of VGG-16 ANN and SNN, obtained empirically from our cycle-accurate simulators, over the CIFAR-10 test set.
Table 15: Comparison of inference energy consumption across different networks.
Type of Network Total Inference Energy (mJ)
ANN (L=4) 3724.865
Default IF (T=4) 445.47
PASC IF (T=4) 446.02
A.10.2 Latency
We would like to emphasize that the proposed PASC IF neuron adds only a minor overhead to inference latency and energy cost, since the end-to-end latency and energy consumption of the network are dominated by the MatMul layers, which incur more operations than the IF layers, as discussed in Section 4.2. We quantify the overhead by implementing the PASC IF neuron on the LoAS accelerator  yin2024loas, which employs a fully temporal-parallel dataflow architecture designed to minimize both data movement across timesteps and the end-to-end latency of sparse SNNs. Table  16 reports the inference latency of VGG-16 SNN, over the CIFAR-10 test set, for both the default IF and PASC IF configurations.
Table 16: Inference latency comparison between different IF types.
Type of IF Total Inference Latency (s)
Default IF (T=4) 0.40
PASC IF (T=4) 0.45
A.11 Comparison between PASCAL and direct SNN training
We compare our method against the direct training approach proposed in  deng2022temporal. The results are presented in Table 17. We show significant improvements in accuracy across datasets, primarily for more complex datasets like ImageNet.
Table 17: Comparison between PASCAL and direct SNN training (TET).
Dataset Method Model Timesteps Accuracy
CIFAR-10 TET ResNet-19 4 94.44
PASCAL (ours) VGG-16 4.21 95.61
CIFAR-100 TET ResNet-19 4 74.47
PASCAL (ours) VGG-16 4.21 76.06
ImageNet TET Spiking ResNet-34 6 64.79
PASCAL (ours) ResNet-34 8.17 74.31